OpenStudy (babynini):

Find exact solutions of the equation in the interval [0,pi) cos y cot^2 y = cos y

OpenStudy (babynini):

Help, please! :)

OpenStudy (babynini):

@Miracrown

OpenStudy (babynini):

Anybody have any ideas? o.o

Nnesha (nnesha):

\[\huge\rm \rm cos y \cot^2 y = \cos y \] divide both side by cos y .....?

OpenStudy (babynini):

so we're left with cot^2y ------ cosy

Nnesha (nnesha):

\[\huge\rm \frac{ \cancel{cos y} \cot^2 y }{\cancel{cos y}}= \frac{\cancel{ \cos y}}{\cancel{cos y}} \] \[\large\rm cot^2 y = 0\]

zepdrix (zepdrix):

Division is badddd +_+ We lose some stuffffff I think subtracting cosy to the left side is probably betterrrrr :D

zepdrix (zepdrix):

\[\Large\rm \cos y \cot^2y-\cos y=0\]:3

Nnesha (nnesha):

=.= okayyyyyyyyyyyyyyy then subtaraarrararaaaact lololol

Nnesha (nnesha):

yeah that's easy ;3

OpenStudy (babynini):

you guys haha ok now what o.0

zepdrix (zepdrix):

nesh, your pictures are better lately 0_o less creepy sad babies = good

zepdrix (zepdrix):

now factor out a cosy from each term :)

OpenStudy (babynini):

which would look like...

OpenStudy (babynini):

cosy (cot^2-1) = 0 ?

zepdrix (zepdrix):

Good c: then apply your `zero factor property`: `cosy=0` and `cot^2y-1=0`

OpenStudy (babynini):

so then the exact solutions are what?

zepdrix (zepdrix):

uhh.. well.. solve for them separately. cos y = 0 what angles does this correspond to? when cosine is giving you 0, which angles? :o

Nnesha (nnesha):

\[\huge\rm (x , y ) \rightarrow (\cos , \sin ) \] ;3

OpenStudy (babynini):

er 0? haha

zepdrix (zepdrix):

oh great the babies are back -_- cosine is 0 at 0 degrees? no.

zepdrix (zepdrix):

cosine of 0 degrees is 1, yes?

zepdrix (zepdrix):

so that's no bueno

OpenStudy (babynini):

pi/2

Nnesha (nnesha):

 lO_Ok at the unit circle o^_^o